OpenStudy (anonymous):

what is the LCD of 5/12 and -9/16

OpenStudy (anonymous):

They both seem to be 1.

OpenStudy (anonymous):

1 isnt an option...

OpenStudy (mathstudent55):

This is how you find the least common multiple of numbers. First, find the prime factors of each number. What are the prime factors of 12? What are the prime factors of 16?

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

cant figure that out help!!!!!!!!!

OpenStudy (mathstudent55):

The prime numbers are: 2, 3, 5, 7, 11, 13, 17, 19, 23, 29, etc. Every whole number, 2 or larger, is either a prime number of a composite number. A copmposite number is a number that is equal to the product of prime numbers.

OpenStudy (mathstudent55):

For example, looking at the list of the first several prime numbers, you see that 3, 5, and 11 are prime. For example, 10 is not on the list. 10 is not prime. 10 is a composite number. It is the product of 2 and 5. 2 and 5 are prime.

OpenStudy (mathstudent55):

What about 12? 12 is not prime. It is 3 * 4, but 4 is not prime, so we can find the complete list of the prime factors of 12. 12 = 2 * 2 * 3, and 2 and 3 are prime.

OpenStudy (anonymous):

so how would i find a prime factor?

OpenStudy (mathstudent55):

Great question. There are several methods. Here is one of the methods. Start with the given number. Divide it by the smallest prime number, which is 2, as many times as you can. This means, no reminder division. Once you can't divide it by that prime number, divide it by the next one. If it is not divisible, skip that prime number and move on to the next prime number. When the division gives 1 as the result you are done. Then make a list of prime factors.

OpenStudy (mathstudent55):

Here is an example. Find the prime factors of 36. 36 is even. Let's divide it by 2. 36/2 = 18 18 is also divisible by 2: 18/2 = 9 9 is not divisible by 2, so we move on to the next prime number, 3. 9/3 = 3 3/3 = 1 Since we arrived at 1, our process is over. Now we see all the prime numbers we divided by: 2, 2, 3, 3 That means 36=2×2×3×3=22×32